OpenStudy (anonymous):

Find the slope of the line. y varies directly with x: y = -2 when x = 4. Find the constant of variation. Then find the value of y when x = -1/2.

OpenStudy (jdoe0001):

\(\begin{array}{cccccclllll} \textit{something}&&\textit{varies directly to}&&\textit{something else}\\ \quad \\ \textit{something}&=&{\color{brown}{ \textit{some value}}}&\cdot &\textit{something else}\\ \quad \\ y&=&{\color{brown}{ n}}&\cdot&x \\\hline\\ && y={\color{brown}{ n }}x \end{array} \\ \quad \\ \bf y=-2\qquad x=4\quad \implies -2={\color{brown}{ n }}4\impliedby find\ {\color{brown}{ n}}\textit{ or constant of variation} \) once you find it, plug it back in the equation and solve when x = -1/2

OpenStudy (anonymous):

@jdoe0001 - the constant of variation is -2. so when x = -1/2, y = 1...?

OpenStudy (jdoe0001):

-2? are you sure?

OpenStudy (jdoe0001):

\(\bf y=-2\qquad x=4\quad \implies -2={\color{brown}{ n }}4\implies \square ?={\color{brown}{ n}}\)

OpenStudy (anonymous):

I'm confused, sorry. @jdoe0001

OpenStudy (jdoe0001):

well... lemme write it a bit differently if you have -2 = 4n what would "n" be?

OpenStudy (anonymous):

I think I did it backwards, sorry. Is it -1/2? @jdoe0001

OpenStudy (jdoe0001):

yeap is -1/2 now we know that n = -1/2 thus what's "y" when x = -1/2? thus \(\bf y=-2\qquad x=4\quad \implies -2={\color{brown}{ n }}4\implies -\cfrac{1}{2}={\color{brown}{ n}} \\ \quad \\ y=-\cfrac{1}{2}x\qquad\qquad x=-\cfrac{1}{2}\implies y=-\cfrac{1}{2}\left( -\cfrac{1}{2} \right) \\ \quad \\ y=-\cfrac{1}{2}\cdot -\cfrac{1}{2}\implies y=\cfrac{-1\cdot -1}{2\cdot 2}\implies y=\cfrac{1}{4}\)

OpenStudy (anonymous):

@jdoe0001 - 1/4

OpenStudy (jdoe0001):

minus * minus = positive

OpenStudy (anonymous):

@jdoe0001 - I'm sorry if I confused you. I tend to put a "-" when I reply to someone. my answer is 1/4.

OpenStudy (jdoe0001):

heeh tis ok =)